OpenStudy (anonymous):

For what value of x is the line tangent to y=x^2 parallel to the line tangent to y=x^.5?

OpenStudy (mathmate):

when f1'(x) = f2'(x) or 2x = 1/(2x) x=1/2 ALternatively, since f1 is the inverse of f2, so when f1'(x)=f2'(x)=1. Comes to the same answer.

OpenStudy (mathmate):